### v3

@@ 3 Quality-based Selection
In this section, we present methods on quality assessment and selection. Without lose of generality, the term “quality" here primarily refers to the integrity, accuracy, and rationality of instruction-response datapoints. For integrity, it measures whether the instruction and response are understandable and complete in both format and content. For accuracy, it estimates whether the “ground-truth" response truly corresponds to the instruction. For rationality, we focus on the consistency and coherency of the instruction context. Although these three dimensions all contribute to the overall quality, in general, existing methods often formulate a unified scoring mechanism to implicitly consider them partially or comprehensively.
@@ 3.1 Hand-crafted Indicators
Traditional methods develop hand-crafted indicators to evaluate the data quality in terms of linguistic analysis such as vocabulary, syntax, and inter-sample semantic similarity. Each indicator is manually, empirically designed with prior knowledge on the language, domain, and task of the corpus under investigation. The calculation of each indicator is explicitly defined and does not require training and inference of proxy models or language models. Although the indicators are hand-crafted, deep learning models such as sentence encoders might be leveraged to extract embedding representations for each instruction text. For the datapoint $x_{i}$, its indicator $IND_{i}$ can be typically defined as:
@@
where $M$ denotes the total number of indicators and $f$ is the aggregation function which depends on both the instruction task and dataset. One can simply use linear combination with pre-defined or dynamically adjusted weights while meticulous tuning might be needed for the ultimate $f$. Given the indicators $IND_{i}$ for each $x_{i}$, two intuitive selection methods can be adopted: 1) to filter out datapoints whose indicator scores are below a pre-defined threshold; 2) to keep only the samples whose indicator scores rank within a certain range of percentiles. Mathematically, these two selection mechanism can be respectively represented as:
@@
Mishra et al. (2020b) and Mishra et al. (2020a) introduce a data quality metric, namely the DQI, to quantify the differences between successive benchmarks by giving high scores to generalizable samples and low scores to biased samples. Such a metric implies whether a well-trained model truly learns the underlying task rather than overfitting the spurious bias of specific benchmarks. Specifically, DQI has seven components including vocabulary, inter-sample N-gram frequency and relation, inter-sample semantic textual similarities (STS), intra-sample word similarity, intra-sample STS, N-Gram frequency per label, and inter-split STS. Based on the proposed DQI, Mishra and Sachdeva (2020) proposes to prune existing huge NLP datasets and demonstrates that the model trained on only  2% of the SNLI dataset achieves near-equal performance with that on the entire set. It first performs AFLite Le Bras et al. (2020), which is detailed in , to keep samples with predictability scores over a pre-defined threshold and then delete bottom $k$ samples with lowest DQI scores. Dang and Verma (2024) further split DQI components into linguistic indicators and semantic indicators, and validate their respective roles in detecting outliers, noises, and duplications. Apart from training-oriented data selection, quality indicators can also be employed to identify the most discriminative samples in the evaluation set to expedite evaluation of LLMs. Saranathan et al. investigates key indicators such as spelling errors Yannakoudakis and Fawthrop (1983), average word length, excessive word repetition, and the compound probability distribution. These indicators stem from the traditional studies on text readability (i.e., readability formulas and sophisticated features) Klare et al. (1963, 1984); Dubay (2004); Kintsch and Vipond (2014); Kemper (1983). Recent studies on readability leverage NLP systems to extract more advanced and informative features for readability measures Si and Callan (2001); Collins-Thompson and Callan (2005); Schwarm and Ostendorf (2005); Feng et al. (2010). François (2010, 2011); François and Fairon (2012) systematically analyze the lexical features, syntactic features, semantic features, and language-specific features with up to 46 indicators. François and Miltsakaki (2012) validates these manually-designed (classical) and NLP-enabled (non-classical) readability formulas, implying that high-quality text corpus can be pinpointed by such carefully designed metrics. Felice and Specia (2012) finds that the hand-crafted linguistic features should be combined with other shallow features for better quality estimation.
@@ 3.2 Model-based Indicators
The model-based indicators, on the other hand, leverage trainable models to predict the indicators for each instruction datapoint. The trainable models used for data quality measurement can either share the same or similar architecture with the language model under development, or possesses completely different implementation choices. Accordingly, these indicators can be simply defined as:
$IND_{i}$ $=f(IND^{1}_{\theta}(x_{i}),IND^{2}_{\theta}(x_{i}),$ (6)
$IND^{3}_{\theta}(x_{i}),...IND^{M}_{\theta}(x_{i})),$
where the learnable parameters $\theta$ highlight the difference between model-based and hand-crafted indicators. Based on the computed indicators, similar selection mechanisms (Eqs. 45) can be adopted to select favorable datapoints.
@@
One of the most intuitive model-based indicators is perplexity Shannon (2001); Jelinek et al. (1977); Jelinek (1980). It is frequently mentioned as the evaluation metric for pre-trained language models Penedo et al. (2023); Radford et al. (2018, 2019); Brown et al. (2020); Achiam et al. (2023) but can also be employed as a data quality indicator. Ankner et al. (2024) proposes to use a small GPT-style reference model such as MPT 125M Team (2023) to prune dataset via perplexity-based sampling for training a 3B model. Specifically, for any datapoint $x_{i}$, the perplexity is defined as the exponential of negative likelihood with base of $2$:
@@
Based on the perplexity inferred from a small model, samples at the high and medium percentiles are chosen by Eq. 5 for downstream fine-tuning.  Deng et al. (2021) develops a unified evaluator framework to score the generated outputs for natural language generation tasks. A RoBERTa-based Liu et al. (2019) discriminator learns to score responses in terms of consistency, relevance, preservation, engagingness, and groundedness. One could simply adopt such a discriminator for evaluation of the instruction-response pairs. Zhong et al. (2022) further proposes a multi-dimensional scoring evaluator. For each evaluation dimension, the original ground-truth instruction-response pairs are converted into positive samples in the form of boolean question-answer problems. The negative samples are respectively constructed via rule-based transformation. The evaluator itself is implemented as T5 model Raffel et al. (2020) and trained on these positive and negative samples for scoring in the range from 0 to 1.  Jiang et al. (2024c) prunes the UltraChat Ding et al. (2023) dataset by scoring each datapoint by learning complexity of a pre-trained Qwen-1.8B model Bai et al. (2023). Specifically, the learning complexity is calculated as the averaged prediction confidence of different subnets:
$\tilde{S}(x_{i})=\frac{1}{I}\sum^{I}_{j=1}PPLX_{i;\Theta_{j}}^{-1},$ (8)
where $I$ is the number of subnets. Each subnet $\Theta_{j}$ is obtained by adjusting the dropout rate from 10% to 90% on the original $\Theta$ of any pre-trained language model. Instruction datapoints with small $\tilde{S}(x_{i})$ are easy ones and should be kept first during pruning. Both  Bukharin and Zhao (2023) and  Du et al. (2023) employ reward models to assess the quality of each instruction pairs. They respectively utilize the raft model Dong et al. (2023) and the deberta-v3-large-v2  [^1] https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2 for reward scoring:
@@
where $r_{\theta}$ denotes the reward model. $t$ is the index where $x_{i(<t)}$ and $x_{i(\geq t)}$ respectively denote the instruction $Q$ and response $A$.  Marion et al. (2023) investigates three classic metrics in clean set selection Guo et al. (2022); Song et al. (2022); Natarajan et al. (2013); Qin et al. (2024): perplexity (Eq. 7), error $l_{2}$-Norm (EL2N) Paul et al. (2021), and memorization ranking Biderman et al. (2024). Specifically, EL2N is defined as:
@@
where $\mathbf{y}_{i(j)}\in\mathbb{R}^{N_{\text{vocab}}}$ denotes the one-hot vector of the vocabulary size $N_{\text{vocab}}$, where its element indexed at $x_{i(j)}$ equals one. The memorization ranking is represented as:
@@
where $N_{\text{win}}$ denotes the length of a consecutive sequence and $M_{\text{offset}}$ is an offset of the starting index. The $\hat{x}_{i(M_{\text{offset}}+j)}$ refers to the generated token given input $x_{i(<{M_{\text{offset}}}+j)}$, and $x_{i(M_{\text{offset}}+j)}$ is its ground-truth. Cao et al. (2023) combines both hand-crafted indicators (e.g., input length, output length, MTLD McCarthy and Jarvis (2010), and kNN-i Dong et al. (2011)) and model-based indicators (e.g., reward score, perplexity, and Uni-Eval metrics Zhong et al. (2022)) for fitting the loss of a LLM on the evaluation set. The linear regression model is optimized via least squares method Bjork (1988) and the optimal selection of instruction data is achieved via BlendSearch Wang et al. (2021a, b) for minimizing the estimated evaluation loss. Li et al. (2023a) is one of the most pioneering works that leverages the target language model itself to perform self-guided data selection. The language model is first “warmed-up" with very few samples randomly chosen from the pool to learn from brief experience. Then, such an experienced model evaluates each instruction-response pair via the instruction-following difficulty (IFD) score. The IFD score measures how much guidance or assistance the instruction provides to the generation of ground-truth response, by comparing the loss of causal language modeling on the response with and without instruction:
@@
where the index $t$ splits apart the instruction $Q$ and the response $A$. Samples whose IFD scores over $\tau_{\text{max}}=1$ are invalid datapoints with misaligned, mismatched instruction-response pairs. The empirical setting of $\tau_{\text{min}}$ affects the trade-off between quality and diversity of the selected datapoints. Zhao and Fang (2024) comprehensively employs hand-crafted indicators for low-level quality filtering, and uses perplexity and IFD score for high-level filtering. A voting mechanism is additionally introduced with IFD scores from one pre-trained base model and one fine-tuned experience model.  Li et al. (2024b) corroborates that both the perplexity and IFD scores inferred from a rather small GPT2-125M Radford et al. (2019) are indicative in selecting high-quality datapoints for training LLaMA2-7B and LLaMA2-13B Touvron et al. (2023b), which greatly improves selection efficiency.
Another popular model-based quality filtering method is AF-Lite Le Bras et al. (2020), which has been applied and validated in recent NLP studies Mishra and Sachdeva (2020); Sakaguchi et al. (2021). It randomly partition all available datapoints into training set and validation set. Then, a model (e.g., linear classifier or language model) is trained on the training set and inferred on the validation set. Such process iterates $m$ times for calculation of the predictability score, which is defined as the ratio of the number of correctly predicted response over the number of total predictions:
@@
Bhatt et al. (2024) presents uncertainty-based quality indicators such as mean entropy Settles (2011); Kremer et al. (2014), least confidence Settles (1995, 2011), mean margin Tong and Koller (2001); Balcan et al. (2006); Settles (2011), and min margin Nguyen et al. (2022). Mathematically, such uncertainty indicators are defined as:
@@
where $\beta_{1}$ and $\beta_{2}$ denote the largest and second largest element of the probability $P(x_{i(<j)};\theta)\in\mathbb{R}^{N_{\text{vocab}}}$ for the newly generated $j$-th token. However,  Wu et al. (2023) finds that such uncertainty-based data sampling methods perform worse than random sampling on Databricks-Dolly Conover et al. (2023), SelfInstruct-Davinci Taori et al. (2023), and SelfInstruct-GPT4 Peng et al. (2023).
@@
Hybrid techniques that simultaneously combines perplexity, uncertainty, reward scores, and other training-aware metrics are promising in selecting unbiased high quality samples. In consideration of the training and inference cost, it is feasible to employ small proxy models as alternatives for computing model-based indicators.

### v4

@@ 3 Quality-based Selection
In this section, we present methods on quality assessment and selection. Without lose of generality, we present the unified formulation of quality measurement. The evaluation function $q(x_{i})$ in quality-based methods can be decomposed into two fine-grained parts: 1) instruction quality $q_{I}(x_{i})$ and 2) response quality $q_{R}(x_{i})$.
$q(x_{i})=f_{q}(q_{I}(x_{i<t}),q_{R}(x_{i\geq t})),$ (5)
where $f_{q}$ is an aggregation function that combines the instruction and response quality scores either explicitly or implicitly. Specifically, the instruction quality $q_{I}$ can be further broken down into: 1) clarity $q_{I}^{C}$ that measures the ease of understanding the task, 2) accuracy $q_{I}^{A}$ that measures how well the instruction aligns with the expected task, and 3) explicitness $q_{I}^{E}$ that measures how explicitly the instruction defines the output constraints (e.g., formats and styles). Consequently, we have $q_{I}(x_{i<t})=g_{I}(q_{I}^{C}(x_{i<t}),q_{I}^{A}(x_{i<t}),q_{I}^{E}(x_{i<t}))$ with the aggregation function $g_{I}$. Similarly for the measurement of response, its quality $q_{R}$ can be assessed via: 1) correctness $q_{R}^{C}$ that measures whether the response correctly answers the instruction, 2) coherence $q_{R}^{H}$ that measures the logical consistency of the response, and 3) pertinence $q_{R}^{P}$ that measures the relevance of the response to the instruction. The final response can be judged as $q_{R}(x_{i\geq t})=g_{R}(q_{R}^{C}(x_{i<t}),q_{R}^{H}(x_{i<t}),q_{R}^{P}(x_{i<t}))$ with an aggregation function $g_{R}$. It is noted that all the mentioned quality measurement components above are only demonstrative and are not enforced explicitly in the development of existing quality-based methods. Therefore, certain fine-grained aspects might be integrated into one formulation without extra considerations.
@@ 3.1 Hand-crafted Indicators
Traditional methods develop hand-crafted indicators to evaluate the data quality in terms of linguistic analysis such as vocabulary, syntax, and inter-sample semantic similarity. Each indicator is manually, empirically designed with prior knowledge on the language, domain, and task of the corpus under investigation. The calculation of each indicator is explicitly defined and does not require training and inference of proxy models or language models. Although the indicators are hand-crafted, deep learning models such as sentence encoders might be leveraged to extract embedding representations for each datapoint $x_{i}$. An indicator $IND_{i}$ can be typically defined as:
@@
where $M$ denotes the total number of indicators and $f$ is the aggregation function which depends on both the instruction task and dataset. One can simply use linear combinations with pre-defined or dynamically adjusted weights. However, meticulous tuning might be needed for the ultimate $f$. Given the indicators $IND_{i}$ for each $x_{i}$, two intuitive selection methods can be adopted: 1) to filter out datapoints whose indicator scores are below a threshold; 2) to keep only the samples whose indicator scores rank within a certain range of percentiles. Mathematically, these two selection mechanisms can be respectively represented as:
@@
Mishra et al. (2020a) and Mishra et al. (2020b) introduce a data quality metric, namely the DQI, to quantify the differences between successive benchmarks by giving high scores to generalizable samples and low scores to biased samples. Such a metric implies whether a well-trained model truly learns the underlying task rather than overfitting the spurious bias of specific benchmarks. Specifically, DQI has seven components including vocabulary, inter-sample N-gram frequency and relation, inter-sample semantic textual similarities (STS), intra-sample word similarity, intra-sample STS, N-Gram frequency per label, and inter-split STS. Based on the DQI, Mishra & Sachdeva (2020) propose to prune existing huge NLP datasets and demonstrates that the model trained on only  2% of the SNLI dataset achieves near-equal performance with that on the entire set. It first performs AFLite Le Bras et al. (2020), which is detailed in Sakaguchi et al. (2021), to keep samples with predictability scores over a threshold and then delete bottom $k$ samples with the lowest DQI scores. Dang & Verma (2024) further split DQI components into linguistic indicators and semantic indicators, and validate their respective roles in detecting outliers, noises, and duplications. Apart from data selection for training LLMs, quality indicators can also be employed to identify the most discriminative samples in the evaluation set to expedite evaluation of LLMs. Saranathan et al. investigate key indicators such as spelling errors Yannakoudakis & Fawthrop (1983), average word length, excessive word repetition, and the compound probability distribution. These indicators stem from the traditional studies on text readability (i.e., readability formulas and sophisticated features) Klare et al. (1963; 1984); Dubay (2004); Kintsch & Vipond (2014); Kemper (1983). Recent studies on readability leverage NLP systems to extract more advanced and informative features for readability measures Si & Callan (2001); Collins-Thompson & Callan (2005); Schwarm & Ostendorf (2005); Feng et al. (2010). François (2010; 2011); François & Fairon (2012) systematically analyze the lexical features, syntactic features, semantic features, and language-specific features with up to 46 indicators. François & Miltsakaki (2012) validate these manually-designed (classical) and NLP-enabled (non-classical) readability formulas, implying that high-quality texts can be pinpointed by such carefully designed metrics. Felice & Specia (2012) find that the hand-crafted linguistic features should be combined with other shallow features for better quality estimation.
@@ 3.2 Model-based Indicators
The model-based indicators, on the other hand, leverage trainable models to predict the indicators for each datapoint. The trainable models used for data quality measurement can either share the same or similar architecture with the language model under development, or possess completely different implementations. Accordingly, these indicators can be simply defined as:
$IND_{i}$ $=f(IND^{1}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\theta_{1}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}(x_{i}),IND^{2}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\theta_{2}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}(x_{i}),$ (9)
$IND^{3}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\theta_{3}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}(x_{i}),...IND^{M}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\theta_{M}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}(x_{i})),$
where the learnable parameters $\theta_{1}$, $\theta_{2}$, …, $\theta_{M}$ highlight the difference between model-based and hand-crafted indicators. Based on such indicators, similar selection mechanisms (Eqs. 7 and 8) can be adopted.
@@
One of the most intuitive model-based indicators is perplexity Shannon (2001); Jelinek et al. (1977); Jelinek (1980). It is frequently mentioned as the evaluation metric for pre-trained language models Penedo et al. (2023); Radford et al. (2018; 2019); Brown et al. (2020); Achiam et al. (2023) but can also be employed as a data quality indicator. Ankner et al. (2024) propose to use a small GPT-style reference model such as MPT 125M Team (2023) to prune datasets via perplexity-based sampling for training a 3B model. For any datapoint $x_{i}$, the perplexity is defined as the exponential of negative likelihood with base of $2$:
@@
Based on the perplexity inferred from a small model, samples at the high and medium percentiles are chosen by Eq. 8 for downstream fine-tuning.  Deng et al. (2021) develop a unified evaluator framework to score the generated outputs for natural language generation tasks. A RoBERTa-based Liu et al. (2019) discriminator learns to score responses in terms of consistency, relevance, preservation, engagingness, and groundedness. One could simply adopt such a discriminator for evaluation of instruction-response pairs. Zhong et al. (2022) further propose a multi-dimensional scoring evaluator. For each evaluation dimension, the original instruction-response pairs are converted into positive samples in the form of boolean question-answer problems. The negative samples are respectively constructed via a rule-based transformation. The evaluator itself is implemented as a T5 model Raffel et al. (2020) and trained on these positive and negative samples for scoring in the range from 0 to 1.  Jiang et al. (2024c) prune the UltraChat Ding et al. (2023) dataset by scoring each datapoint by the learning complexity of a pre-trained Qwen-1.8B model Bai et al. (2023). Specifically, the learning complexity is calculated as the averaged prediction confidence of different subnets:
$\tilde{S}(x_{i})=\frac{1}{I}\sum^{I}_{j=1}PPLX_{i;\Theta_{j}},$ (11)
where $I$ is the number of subnets. Each subnet $\Theta_{j}$ is obtained by adjusting the dropout rate from 10% to 90% on the original $\Theta$ of any pre-trained language model. Such a dropout technique imitates the learning progress in a training-free manner. Datapoints with small $\tilde{S}(x_{i})$ are considered as easy ones. They are often learned at an earlier training stage with less model capacity compared with hard ones. In a data-poor regime, easy datapoints are more informative and should be kept first. On the contrary, in a data-rich regime, hard datapoints should be treasured in that more samples can be preserved for fine-tuning. Both  Bukharin & Zhao (2023) and  Du et al. (2023) employ reward models to assess the quality of each instruction pairs. They respectively utilize the raft model Dong et al. (2023) and the deberta-v3-large-v2  [^1] https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2 for reward scoring:
@@
where $r_{\theta}$ denotes the reward model. $t$ is the index where $x_{i(<t)}$ and $x_{i(\geq t)}$ respectively denote the instruction $Q$ and response $A$.  Marion et al. (2023) investigate three classic metrics in clean set selection Guo et al. (2022); Song et al. (2022); Natarajan et al. (2013); Qin et al. (2024): perplexity (Eq. 10), error $l_{2}$-Norm (EL2N) Paul et al. (2021), and memorization ranking Biderman et al. (2024). Specifically, EL2N is defined as:
@@
where $\mathbf{y}_{i(j)}$ denotes the one-hot ground-truth vector as the target of the probability vector $P(x_{i(<j)};\theta)$. Specifically, both $P(x_{i(<j)};\theta)\in\mathbb{R}^{N_{\text{vocab}}}$ and $\mathbf{y}_{i(j)}\in\mathbb{R}^{N_{\text{vocab}}}$ are of the same dimension, where $N_{\text{vocab}}$ denotes the vocabulary size associated with the tokenizer. For the vector of $\mathbf{y}_{i(j)}$, all elements are zero except that the element indexed at $x_{i(j)}$ is one. The memorization ranking is represented as:
@@
where $N_{\text{win}}$ denotes the length of a consecutive sequence and $M_{\text{offset}}$ is an offset of the starting index. The $\hat{x}_{i(M_{\text{offset}}+j)}$ refers to the generated token given input $x_{i(<{M_{\text{offset}}}+j)}$, and $x_{i(M_{\text{offset}}+j)}$ is its ground-truth. Cao et al. (2023) combine both hand-crafted indicators (e.g., input length, output length, MTLD McCarthy & Jarvis (2010), and kNN-i Dong et al. (2011)) and model-based indicators (e.g., reward score, perplexity, and Uni-Eval metrics Zhong et al. (2022)) for fitting the loss of a LLM on the evaluation set. The linear regression model is optimized via least squares method Bjork (1988) and the optimal selection of instruction data is achieved via BlendSearch Wang et al. (2021a; b) for minimizing the estimated evaluation loss. Li et al. (2023a) propose one of the most pioneering works that leverages the target language model itself to perform self-guided data selection. The language model is first "warmed-up" with very few samples randomly chosen from the pool to learn from brief experience. Then, such an experienced model evaluates each instruction-response pair via the instruction-following difficulty (IFD) score. The IFD score measures how much guidance or assistance the instruction provides to the generation of ground-truth response, by comparing the loss of causal language modeling on the response with and without instruction:
@@
where the index $t$ splits apart the instruction $Q$ and the response $A$. Samples whose IFD scores over $\tau_{\text{max}}=1$ are invalid datapoints with misaligned, mismatched instruction-response pairs. The empirical setting of $\tau_{\text{min}}$ affects the trade-off between quality and diversity of the selected datapoints. Zhao & Fang (2024) comprehensively employ hand-crafted indicators for low-level quality filtering, and uses perplexity and IFD score for high-level filtering. A voting mechanism is additionally introduced with IFD scores from one pre-trained base model and one fine-tuned experience model.  Li et al. (2024b) corroborate that both the perplexity and IFD scores inferred from a rather small GPT2-125M Radford et al. (2019) are indicative in selecting high-quality datapoints for training LLaMA2-7B and LLaMA2-13B Touvron et al. (2023b), which greatly improves selection efficiency.
Another popular model-based quality filtering method is AF-Lite Le Bras et al. (2020), which has been applied and validated in recent NLP studies Mishra & Sachdeva (2020); Sakaguchi et al. (2021). It randomly partition all available datapoints into the training set and the validation set. Then, a model (e.g., linear classifier or language model) is trained on the training set and inferred on the validation set. Such process iterates $m$ times for calculation of the predictability score, which is defined as the ratio of the number of correctly predicted response over the number of total predictions:
@@
Bhatt et al. (2024) present uncertainty-based indicators such as mean entropy Settles (2011); Kremer et al. (2014), least confidence Settles (1995; 2011), mean margin Tong & Koller (2001); Balcan et al. (2006); Settles (2011), and min margin Nguyen et al. (2022). Mathematically, such uncertainty indicators are defined as:
@@
where $\beta_{1}$ and $\beta_{2}$ denote the largest and second largest elements of the probability $P(x_{i(<j)};\theta)\in\mathbb{R}^{N_{\text{vocab}}}$ for the newly generated $j$-th token. However,  Wu et al. (2023) find that such uncertainty-based data sampling methods perform worse than random sampling on Databricks-Dolly Conover et al. (2023), SelfInstruct-Davinci Taori et al. (2023), and SelfInstruct-GPT4 Peng et al. (2023).
@@
Hybrid techniques that simultaneously combine perplexity, uncertainty, reward scores, and other training-aware metrics are promising in selecting unbiased high quality samples. In consideration of the training and inference cost, it is feasible to employ small proxy models as alternatives for computing model-based indicators.